Mary's age is twice Rick's age and one-third of Steve's age. If the sum of their age is 99, how old are they?
Answer by edjones(8007)

m=mary's age
.
m+m/2+3m=99
2m+m+6m=198 mult each side times 2.
9m=198
m=22 divide both sides by 9.
m/2=11 rick's age
3m=66 steve's age
.
Ed